If Leah is 6 years older than Sue, and John is 5 years older than Leah, and the total of their ages is 41. Then how old is Sue?
I know the answer is Sue is 8 - but I need to know how it was solved?

Answer by user_dude2008(1825)

L=lea
J=john
S=sue

"If Leah is 6 years older than Sue" ---> L=S+6 --> S=L-6
" John is 5 years older than Leah" --> J=L+5
"total of their ages is 41" --> L+J+S=41


L+J+S=41
L+(L+5)+(L-6)=41
3L-1=41
3L=42
L=14 <---- Leah is 14
S=L-6=14-6=8 <---- Sue is 8